# LangForce: Bayesian Decomposition of Vision Language Action Models via Latent Action Queries

### v4

@@ Abstract
Vision-Language-Action (VLA) models have shown promise in robot manipulation but often struggle to generalize to new instructions or complex multi-task scenarios. We identify a critical pathology in current training paradigms where goal-driven data collection creates a dataset bias. In such datasets, language instructions are highly predictable from visual observations alone, causing the conditional mutual information between instructions and actions to vanish, a phenomenon we term Information Collapse. Consequently, models degenerate into vision-only policies that ignore language constraints and fail in out-of-distribution (OOD) settings. To address this, we propose LangForce, a novel framework that enforces instruction following via Bayesian decomposition. By introducing learnable Latent Action Queries, we construct a dual-branch architecture to estimate both a vision-only prior $p(a\mid v)$ and a language-conditioned posterior $\pi(a\mid v,\ell)$. We then optimize the policy to maximize the conditional Pointwise Mutual Information (PMI) between actions and instructions. This objective effectively penalizes the vision shortcut and rewards actions that explicitly explain the language command. Without requiring new data, LangForce significantly improves generalization. Extensive experiments across on SimplerEnv and RoboCasa demonstrate substantial gains, including an 8.8% improvement on the challenging OOD SimplerEnv benchmark, validating the ability of our approach to robustly ground language in action.
@@ 1 Introduction
    We demonstrate that LangForce achieves state-of-the-art performance on SimplerEnv and RoboCasa, with a remarkable 8.8% improvement in OOD generalization on SimplerEnv, proving its effectiveness in breaking the vision shortcut.
2 Motivation: The Illusion of Instruction Following
@@ 2.1 Experiment 1: The Vision Shortcut in ID Testing
We first train a standard VLA model on a subset of the Humanoid robot tabletop manipulation data from PhysicalAI-Robotics-GR00T-X-Embodiment-Sim (Bjorck et al., 2025b) and evaluate on 24 tasks from the RoboCasa benchmark (Nasiriany et al., 2024). Averaged across all 24 tasks, the vision-only model achieves a success rate of 44.6%, which is close to the language-conditioned baseline of 47.8%. This small gap reveals that the model can succeed without relying on language instructions, as the training and evaluation scenes and tasks are highly similar, enabling the model to learn a near-deterministic mapping from vision to action. Figure 1 provides a relevant example.
@@ 2.2 Experiment 2: Failure in Ambiguous Scenarios
To further investigate this behavior, we train a standard VLA model on the LIBERO benchmark (Liu et al., 2023), which contains four subsets: Spatial, Object, Long, and Goal. We train on all four training sets and evaluate on all four test sets. The vision-only model achieves success rates comparable to the full VLA model on three subsets (Spatial: 90.2%, Object: 99.6%, Long: 86.0% in Vision-Only, Spatial: 98.8%, Object: 99.0%, Long: 94.4% in Baseline), where each visual scene corresponds to a single task. However, on the LIBERO Goal subset, the vision-only success rate plummets to 9.8% (98.0% in Baseline).
@@
Figure 2: (a) In LIBERO Goal (Liu et al., 2023), the same scene corresponds to multiple tasks, revealing the ambiguity that vision-only models fail to resolve. (b) Action loss curves on BridgeDataV2 (Walke et al., 2023) show that the vision-only model achieves comparable training loss to the full vision-language model, indicating the presence of visual shortcuts even in diverse, in-the-wild datasets.
2.3 Experiment 3: Catastrophic Failure in OOD Generalization
Finally, we test the generalization capability by training on the high-quality BridgeDataV2 dataset (Walke et al., 2023) (diverse, in-the-wild scenes) and evaluating on SimplerEnv (Li et al., 2024d) (simulation, OOD). During training on BridgeDataV2, the vision-only model achieves an action loss of 0.13, comparable to the full language-conditioned model’s loss of 0.08 (as shown in Figure 2(b)). This indicates that even in diverse, in-the-wild scenarios, the model can still identify visual shortcuts (e.g., specific lighting or background features mapping to specific actions) to minimize the training objective without truly grounding the language instructions.
@@ 2.4 Theoretical Insight: Information Collapse
In goal-driven datasets, the deterministic mapping $v\to\ell$ implies $H(\ell\mid v)\approx 0$ (Xu et al., 2025). Consequently, $I(\ell;a\mid v)$ is forced to vanish, theoretically preventing the model from learning any dependency between $a$ and $\ell$ beyond what is already captured by $v$. To break this deadlock, we cannot rely on standard likelihood maximization. Instead, we must explicitly intervene to maximize the information gain provided by the action. This motivates our use of the Log-Likelihood Ratio (LLR), which effectively estimates the Pointwise Mutual Information (PMI), rewarding the policy only when it captures the specific semantics of $\ell$ that are not predictable from $v$.
@@ 3.2 Latent Action Queries
To efficiently instantiate the proposed Bayesian decomposition within a unified VLM architecture, we introduce Latent Action Queries. We extend the VLM vocabulary with $K=64$ learnable tokens, denoted as $\mathcal{Q}=\{\texttt{<|action\_0|>},\dots,\texttt{<|action\_K|>}\}$.
These queries function as a dedicated bottleneck interface between the VLM (e.g., Qwen3-VL (Bai et al., 2025)) and the continuous action head (a Diffusion Transformer (Peebles and Xie, 2023)). Unlike recent VLA architectures such as $\pi_{0}$ (Black et al., 2024) and GR00T (Bjorck et al., 2025b; GEAR-Team et al., 2025), which typically feed the hidden states of all input tokens to the action expert, we append $\mathcal{Q}$ to the input sequence and exclusively use their corresponding hidden states $\mathbf{H}_{\mathcal{Q}}\in\mathbb{R}^{K\times D}$ to condition the action head. This design choice is critical: by leveraging the causal masking inherent in decoder-only VLMs, we can precisely control the information encoded in $\mathbf{H}_{\mathcal{Q}}$ simply by changing the position of $\mathcal{Q}$ in the input sequence. This flexibility enables the strict separation of vision-only and vision-language contexts required for our dual-branch strategy.
@@ 3.3 Dual-Branch Training Framework
Due to the causal attention mask of the decoder-only VLM, the tokens in $\mathcal{Q}$ can attend to the visual observation $v$ but cannot attend to the language instruction $\ell$ (which appears later). Thus, the hidden states $\mathbf{H}_{\mathcal{Q}}^{\text{prior}}$ encode purely visual information. We use these features to predict the action $a$, optimizing a flow-matching loss $\mathcal{L}_{\text{prior}}$ to learn the dataset’s inherent action bias.
@@
In addition to action prediction, we explicitly optimize the LLR objective. We treat the VLM’s language modeling loss as a proxy for $\log p(\ell\mid\dots)$. Specifically, in the Priori Branch, the language tokens $\ell$ attend to $v,\mathcal{Q}]$. Since $\mathcal{Q}$ encodes the action information $a$ (via the prior), the probability of generating $\ell$ in this branch approximates $p(\ell\mid v,a_{\text{prior}})$. In the Posteriori Branch, we can compute a baseline $p(\ell\mid v)$ by detaching gradients or using a separate pass. However, a more direct and numerically stable approach is to maximize the difference in log-probabilities of the language tokens between the two branches. We define the LLR loss as:
@@ 3.4 Total Training Objective
where $\lambda$ balances the contribution of the prior and posterior action losses, and $\beta$ controls the strength of the LLR regularization. During inference, we exclusively execute the Posteriori Branch to obtain $\mathbf{H}_{\mathcal{Q}}^{\text{post}}$ and generate actions via the DiT. This ensures that our method incurs no additional computational overhead compared to standard VLA baselines at test time.
@@ 4 Experiment
To comprehensively evaluate the effectiveness of LangForce, we conducted extensive experiments on two simulation benchmarks: SimplerEnv, and RoboCasa. Our training pipeline is built upon the StarVLA framework (starVLA, 2025), distributed across 16 NVIDIA H100 GPUs, and strictly follows its default training protocols to ensure fair comparison. In our experiments, LangForce is instantiated on the QwenGR00T architecture from StarVLA. Broader simulation benchmarks and real-world robot experiments are currently underway. We employ the AdamW optimizer (Loshchilov and Hutter, 2017) initialized with a learning rate of 1e-5 and a cosine annealing schedule. System-level optimizations include DeepSpeed ZeRO-2 (Rasley et al., 2020), gradient clipping at a norm of 1.0, and no gradient accumulation. All baseline performance metrics are obtained from their original papers or other peer-reviewed publications.
@@
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA (Kim et al., 2024) 4.2 0.0 0.0 12.5 4.2 OpenVLA-OFT (Kim et al., 2025) 12.5 4.2 4.2 72.5 23.4 RoboVLM (Li et al., 2024c) 50.0 37.5 0.0 83.3 42.7 CogACT (Li et al., 2024b) 71.7 50.8 15.0 67.5 51.3 SpatialVLA (Qu et al., 2025) 20.8 20.8 25.0 70.8 34.4 TraceVLA (Zheng et al., 2025b) 12.5 16.6 16.6 65.0 27.7 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.2 62.5 29.2 91.6 53.1 $\pi_{0.5}$ (Black et al., 2025) 49.3 64.7 44.7 69.7 57.1 Isaac-GR00T-N1.6-Bridge (GEAR-Team et al., 2025) 64.5 65.5 5.5 93.0 57.1 QwenGR00T (Baseline) + Qwen3-VL-4B (starVLA, 2025) 87.5 50.0 29.2 64.2 57.7 LangForce + Qwen3-VL-4B 89.6 63.8 33.3 79.2 66.5
4.1 Experiments on SimplerEnv
We utilize two large-scale subsets from the Open X-Embodiment (OXE) dataset: BridgeDataV2 (Walke et al., 2023) and Fractal (Brohan et al., 2022). The model is fine-tuned for 40k steps on a cluster of 16 GPUs (batch size 16 per device). This benchmark includes four manipulation tasks: “Put spoon on towel”, “Put carrot on plate”, “Stack green cube on yellow cube”, and “Put eggplant in yellow basket”. For each task, we evaluate the VLA policies using the official evaluation scripts provided by the SimplerEnv repository (Li et al., 2024d). To mitigate the effects of randomness, we run 480 independent trials and report the average performance (Avg@480).
Table 2: Results of evaluating the VLA models with the GR1 robot in the RoboCasa Tabletop simulation environment. The results for the first five baseline methods are sourced from the official starVLA experiments (starVLA, 2025). We highlight the best results in bold and the second-best results with underline.
Task Isaac-GR00T N1.6 QwenGR00T +Qwen3VL QwenPI +Qwen3VL QwenOFT +Qwen3VL QwenFAST +Qwen3VL VisionOnly QwenGR00T LangForce +Qwen3VL
PnP Bottle To Cabinet Close 51.5 46.0 26.0 30.0 38.0 70.0 60.0
PnP Can To Drawer Close 13.0 80.0 62.0 76.0 44.0 78.0 72.0
PnP Cup To Drawer Close 8.5 54.0 42.0 44.0 56.0 42.0 46.0
PnP Milk To Microwave Close 14.0 48.0 50.0 44.0 44.0 50.0 54.0
PnP Potato To Microwave Close 41.5 28.0 42.0 32.0 14.0 44.0 32.0
PnP Wine To Cabinet Close 16.5 46.0 32.0 36.0 14.0 40.0 56.0
PnP Novel From Cuttingboard To Basket 58.0 48.0 40.0 50.0 54.0 58.0 62.0
PnP Novel From Cuttingboard To Cardboardbox 46.5 40.0 46.0 40.0 42.0 26.0 44.0
PnP Novel From Cuttingboard To Pan 68.5 68.0 60.0 70.0 58.0 72.0 68.0
PnP Novel From Cuttingboard To Pot 65.0 52.0 40.0 54.0 58.0 50.0 40.0
PnP Novel From Cuttingboard To Tieredbasket 46.5 56.0 44.0 38.0 40.0 20.0 36.0
PnP Novel From Placemat To Basket 58.5 42.0 44.0 32.0 36.0 48.0 46.0
PnP Novel From Placemat To Bowl 57.5 44.0 52.0 58.0 38.0 32.0 42.0
PnP Novel From Placemat To Plate 63.0 48.0 50.0 52.0 42.0 34.0 70.0
PnP Novel From Placemat To Tieredshelf 28.5 18.0 28.0 24.0 18.0 16.0 24.0
PnP Novel From Plate To Bowl 57.0 60.0 52.0 60.0 52.0 26.0 34.0
PnP Novel From Plate To Cardboardbox 43.5 50.0 40.0 50.0 30.0 38.0 48.0
PnP Novel From Plate To Pan 51.0 54.0 36.0 66.0 48.0 44.0 52.0
PnP Novel From Plate To Plate 78.7 70.0 48.0 68.0 50.0 60.0 70.0
PnP Novel From Tray To Cardboardbox 51.5 38.0 34.0 44.0 28.0 50.0 60.0
PnP Novel From Tray To Plate 71.0 56.0 64.0 56.0 34.0 64.0 66.0
PnP Novel From Tray To Pot 64.5 50.0 44.0 62.0 46.0 52.0 62.0
PnP Novel From Tray To Tieredbasket 57.0 36.0 50.0 54.0 36.0 42.0 44.0
PnP Novel From Tray To Tieredshelf 31.5 16.0 28.0 30.0 16.0 16.0 22.0
Average 47.6 47.8 43.9 48.8 39.0 44.7 50.4
@@ 4.1 Experiments on SimplerEnv
The results are summarized in Table 1. LangForce consistently outperforms comparison baselines, achieving a state-of-the-art average success rate of 66.5%. Notably, compared to the direct baseline QwenGR00T (55.2%) built on the same StarVLA framework, our method delivers an absolute improvement of 8.8%, validating that the performance gain stems from our proposed Bayesian decomposition rather than the base architecture. Significant improvements are observed in tasks requiring precise object identification and manipulation, such as “Put Carrot on Plate” (+13.6%) and “Put Eggplant in Yellow Basket” (+15.0%). Furthermore, LangForce surpasses other recent strong competitors, including the flow-matching-based $\pi_{0.5}$ (57.1%) and the dual-system Isaac-GR00T-N1.6 (57.1%). These results confirm that by explicitly optimizing the mutual information between language and action, LangForce effectively mitigates the vision shortcut. Fundamentally, this validates that our approach prevents the policy from collapsing into a spurious vision-only prior $p(a|v)$ caused by dataset determinism, and instead compels the model to learn the true causal dependency of actions on language instructions.

### v5

@@ Abstract
Vision-Language-Action (VLA) models have shown promise in robot manipulation but often struggle to generalize to new instructions or complex multi-task scenarios. We identify a critical pathology in current training paradigms where goal-driven data collection creates a dataset bias. In such datasets, language instructions are highly predictable from visual observations alone, causing the conditional mutual information between instructions and actions to vanish, a phenomenon we term Information Collapse. Consequently, models degenerate into vision-only policies that ignore language constraints and fail in out-of-distribution (OOD) settings. To address this, we propose LangForce, a novel framework that enforces instruction following via Bayesian decomposition. By introducing learnable Latent Action Queries, we construct a dual-branch architecture to estimate both a vision-only prior $p(a\mid v)$ and a language-conditioned posterior $\pi(a\mid v,\ell)$. We then optimize the policy to maximize the conditional Pointwise Mutual Information (PMI) between actions and instructions. This objective effectively penalizes the vision shortcut and rewards actions that explicitly explain the language command. Without requiring new data, LangForce significantly improves generalization. Extensive experiments across on SimplerEnv and RoboCasa demonstrate substantial gains, including an 11.3% improvement on the challenging OOD SimplerEnv benchmark, validating the ability of our approach to robustly ground language in action.
@@
footnotetext: Work done at Zhongguancun Academy (Beijing).
@@ 1 Introduction
    We demonstrate that LangForce achieves state-of-the-art performance on SimplerEnv, LIBERO, and RoboCasa, with a remarkable 11.3% improvement in OOD generalization on SimplerEnv, proving its effectiveness in breaking the vision shortcut.
2 Motivation: Vision Shortcut
@@ 2.1 Experiment 1: The Vision Shortcut in ID Testing
We first train a standard VLA model on a subset of the Humanoid robot tabletop manipulation data from PhysicalAI-Robotics-GR00T-X-Embodiment-Sim (Bjorck et al., 2025b) and evaluate on 24 tasks from the RoboCasa benchmark (Nasiriany et al., 2024). Averaged across all 24 tasks, the vision-only model achieves a success rate of 44.6%, which is close to the language-conditioned baseline of 47.8% (as shown in Table 4). This small gap reveals that the model can succeed without relying on language instructions, as the training and evaluation scenes and tasks are highly similar, enabling the model to learn a near-deterministic mapping from vision to action. Figure 1 provides a relevant example.
@@ 2.2 Experiment 2: Failure in Ambiguous Scenarios
To further investigate this behavior, we train a standard VLA model on the LIBERO benchmark (Liu et al., 2023), which contains four subsets: Spatial, Object, Long, and Goal. We train on all four training sets and evaluate on all four test sets. The vision-only model achieves success rates comparable to the full VLA model on three subsets (Spatial: 90.2%, Object: 99.6%, Long: 86.0% in Vision-Only, Spatial: 97.8%, Object: 98.8%, Long: 92.0% in Baseline), where each visual scene corresponds to a single task. However, on the LIBERO Goal subset, the vision-only success rate plummets to 9.8% (97.4% in Baseline).
Figure 2: (a) In LIBERO Goal (Liu et al., 2023), the same scene corresponds to multiple tasks, revealing ambiguity that vision-only models fail to resolve. (b) Action loss curves on BridgeDataV2 (Walke et al., 2023) and Fractal (Brohan et al., 2022) show the vision-only model achieves comparable loss to the full model, indicating the presence of visual shortcuts even in diverse, in-the-wild datasets.
@@
2.3 Experiment 3: Failure in OOD Generalization
Finally, we test the generalization capability by training on the high-quality BridgeDataV2 dataset (Walke et al., 2023) and Fractal (Brohan et al., 2022) dataset (diverse, in-the-wild scenes) and evaluating on SimplerEnv (Li et al., 2024d) (simulation, OOD). During training on BridgeDataV2, the vision-only model achieves an action loss of 0.13, comparable to the full language-conditioned model’s loss of 0.08 (as shown in Figure 2(b)). This indicates that even in diverse, in-the-wild scenarios, the model can still identify visual shortcuts (e.g., specific lighting or background features mapping to specific actions) to minimize the training objective without truly grounding the language instructions.
@@ 2.4 Theoretical Insight: Information Collapse
In goal-driven datasets, the deterministic mapping $v\to\ell$ implies $H(\ell\mid v)\approx 0$ (Xu et al., 2025). Consequently, $I(\ell;a\mid v)$ is forced to vanish, theoretically preventing the model from learning any dependency between $a$ and $\ell$ beyond what is already captured by $v$. To break this deadlock, we cannot rely on standard imitation learning. Instead, we must explicitly intervene to maximize the information gain provided by the action. This motivates our use of the Log-Likelihood Ratio (LLR), which effectively estimates the Pointwise Mutual Information (PMI), rewarding the policy only when it captures the specific semantics of $\ell$ that are not predictable from $v$.
@@ 3.2 Latent Action Queries
To efficiently instantiate the proposed Bayesian decomposition within a unified VLM architecture, we introduce Latent Action Queries. We extend the VLM vocabulary with $K=64$ learnable tokens, denoted as $\mathcal{Q}=\{\texttt{<|action\_1|>},\dots,\texttt{<|action\_K|>}\}$. Correspondingly, the VLM’s embedding layer is expanded with $K$ learnable embedding vectors. These vectors are learned to aggregate task-relevant information from the preceding vision and language tokens into a compact latent representation for action execution.
These queries function as a dedicated bottleneck at the interface between the VLM (e.g., Qwen3-VL (Bai et al., 2025)) and the continuous action head (a Diffusion Transformer (Peebles and Xie, 2023)). Crucially, while the VLM typically processes the full sequence, the bottleneck is enforced by exclusively forwarding the hidden states of these query tokens, $\mathbf{H}_{\mathcal{Q}}\in\mathbb{R}^{K\times D}$, to the action expert. This contrasts with recent VLA architectures such as $\pi_{0}$ (Black et al., 2024) and GR00T (Bjorck et al., 2025b; GEAR-Team et al., 2025), which typically feed the hidden states of all input tokens to the action head.
This design choice is critical: by leveraging the causal masking inherent in decoder-only VLMs, we can precisely control the information encoded in $\mathbf{H}_{\mathcal{Q}}$ simply by changing the position of $\mathcal{Q}$ in the input sequence. This flexibility enables the strict separation of vision-only and vision-language contexts required for our dual-branch strategy.
@@ 3.3 Dual-Branch Training Framework
Due to the causal attention mask of the decoder-only VLM, the tokens in $\mathcal{Q}$ can attend to the visual observation $v$ but cannot attend to the language instruction $\ell$ (which appears later). Thus, the hidden states $\mathbf{H}_{\mathcal{Q}}^{\text{prior}}$ encode purely visual information. We use these features to predict the action $a$, optimizing a flow-matching loss $\mathcal{L}_{\text{prior}}$ to learn the dataset’s inherent action bias. Crucially, we detach $\mathbf{H}_{\mathcal{Q}}^{\text{prior}}$ from the computation graph when optimizing $\mathcal{L}_{\text{prior}}$ to ensure that the gradient updates for the vision-only prior are confined to the DiT action head, preventing the shared VLM backbone from learning visual shortcuts.
@@
In addition to action prediction, we explicitly optimize the LLR objective. We treat the VLM’s language modeling loss as a proxy for $\log p(\ell\mid\dots)$. Specifically, in the Priori Branch, the language tokens $\ell$ attend to $v,\mathcal{Q}]$. Since $\mathcal{Q}$ encodes the action information $a$ (via the prior), the probability of generating $\ell$ in this branch approximates $p(\ell\mid v,a_{\text{prior}})$. In the Posteriori Branch, we can compute a baseline $p(\ell\mid v)$ by detaching gradients. However, a more direct and numerically stable approach is to maximize the difference in log-probabilities of the language tokens between the two branches. We define the LLR loss as:
@@ 3.4 Total Training Objective
where $\lambda$ balances the contribution of the prior and posterior action losses, and $\beta$ controls the strength of the LLR regularization. We set $\lambda=0.3$ and $\beta=0.1$ in our experiments. During inference, we exclusively execute the Posteriori Branch to obtain $\mathbf{H}_{\mathcal{Q}}^{\text{post}}$ and generate actions via the DiT. This ensures that our method incurs no additional computational overhead compared to standard VLA baselines at test time.
@@ 4 Experiment
To comprehensively evaluate the effectiveness of LangForce, we conducted extensive experiments on three simulation benchmarks: SimplerEnv, LIBERO, and RoboCasa. Our training pipeline is built upon the StarVLA framework (starVLA, 2025), distributed across 8 NVIDIA H100 GPUs, and strictly follows its default training protocols to ensure fair comparison. In our experiments, LangForce is instantiated on the QwenGR00T architecture from StarVLA. We employ the AdamW optimizer (Loshchilov and Hutter, 2017) initialized with a learning rate of 1e-5 and a cosine annealing schedule. System-level optimizations include DeepSpeed ZeRO-2 (Rasley et al., 2020), gradient clipping at a norm of 1.0, and no gradient accumulation. All baseline performance metrics are obtained from their original papers or other peer-reviewed publications. To ensure a fair comparison, the training datasets for these baselines encompass the data used in our experiments.
4.1 Experiments on SimplerEnv
We utilize two large-scale subsets from the Open X-Embodiment (OXE) dataset: BridgeDataV2 (Walke et al., 2023) and Fractal (Brohan et al., 2022). The model is fine-tuned for 50k steps on a cluster of 8 GPUs (batch size 16 per device). This benchmark includes four manipulation tasks: “Put spoon on towel”, “Put carrot on plate”, “Stack green cube on yellow cube”, and “Put eggplant in yellow basket”. For each task, we evaluate the VLA policies using the official evaluation scripts provided by the SimplerEnv repository (Li et al., 2024d). To mitigate the effects of randomness, we run 480 independent trials and report the average performance (Avg@480).
@@ 4.1 Experiments on SimplerEnv
Method Put Spoon on Towel Put Carrot on Plate Stack Green Block on Yellow Block Put Eggplant in Yellow Basket Average RT-1-X (O’Neill et al., 2024) 0.0 4.2 0.0 0.0 1.1 Octo-Base (Team et al., 2024) 15.8 12.5 0.0 41.7 17.5 Octo-Small (Team et al., 2024) 41.7 8.2 0.0 56.7 26.7 OpenVLA-OFT (Kim et al., 2025) 34.2 30.0 30.0 72.5 41.8 RoboVLM (Li et al., 2024c) 50.0 37.5 0.0 83.3 42.7 Magma (Yang et al., 2025a) 37.5 29.2 20.8 91.7 44.8 CogACT (Li et al., 2024b) 71.7 50.8 15.0 67.5 51.3 SpatialVLA (Qu et al., 2025) 20.8 20.8 25.0 70.8 34.4 TraceVLA (Zheng et al., 2025b) 12.5 16.6 16.6 65.0 27.7 VideoVLA (Shen et al., 2025) 75.0 20.8 45.8 70.8 53.1 $\pi_{0}$ (Black et al., 2024) 29.2 62.5 29.2 91.6 53.1 $\pi_{0.5}$ (Black et al., 2025) 49.3 64.7 44.7 69.7 57.1 Isaac-GR00T-N1.6-Bridge (GEAR-Team et al., 2025) 64.5 65.5 5.5 93.0 57.1 QwenGR00T (Baseline) + Qwen3-VL-4B (starVLA, 2025) 87.5 50.0 29.2 54.2 55.2 LangForce + Qwen3-VL-4B 89.6 63.8 33.3 79.2 66.5
@@
The results are summarized in Table 1. LangForce consistently outperforms comparison baselines, achieving a state-of-the-art average success rate of 66.5%. Notably, compared to the direct baseline QwenGR00T (55.2%) built on the same StarVLA framework, our method delivers an absolute improvement of 11.3%, validating that the performance gain stems from our proposed Bayesian decomposition rather than the base architecture. Significant improvements are observed in tasks requiring precise object identification and manipulation, such as “Put Carrot on Plate” (+13.6%) and “Put Eggplant in Yellow Basket” (+15.0%). Furthermore, LangForce surpasses other recent strong competitors, including the flow-matching-based $\pi_{0.5}$ (57.1%) and the dual-system Isaac-GR00T-N1.6 (57.1%). These results confirm that by explicitly optimizing the mutual information between language and action, LangForce effectively mitigates the vision shortcut. Fundamentally, this validates that our approach prevents the policy from collapsing into a spurious vision-only prior $p(a|v)$ caused by dataset determinism, and instead compels the model to learn the true causal dependency of actions on language instructions.
4.2 Experiments on LIBERO
Table 2: Comparison on the LIBERO benchmark. We train one policy for all 4 suites. Avg@500 success rates (%) across four task suites: Spatial, Object, Goal, and Long.
Method Spatial Object Goal Long Avg
OpenVLA (Kim et al., 2024) 87.4 88.4 79.2 53.7 76.5
OpenVLA-OFT (Kim et al., 2025) 97.6 98.4 97.9 94.5 97.1
$\pi_{0}$ (Black et al., 2024) 96.8 98.8 95.8 85.2 94.1
$\pi_{0.5}$ (Black et al., 2025) 98.8 98.2 98.0 92.4 96.9
Qwen3-VL-FAST 97.3 97.4 96.3 90.6 95.4
Qwen3-VL-OFT 97.8 98.6 96.2 93.8 96.6
Qwen3-VL-GR00T 97.8 98.8 97.4 92.0 96.5
VisionOnly Qwen3-VL-GR00T 90.2 99.6 9.8 86.0 71.4
LangForce 99.2 99.6 99.4 95.2 98.4
We also evaluate LangForce on the LIBERO benchmark (Liu et al., 2023). Given that the training and testing environments in LIBERO are highly similar and current VLA research has largely saturated this benchmark (with baselines exceeding 95%), our method yields comparable performance on the Spatial, Object, and Long suites. However, on the Goal suite, LangForce achieves a success rate of 99.4%, outperforming the Qwen3-VL-GR00T baseline (97.4%) by +2.0%. As highlighted in Section 2.2, the Goal suite features significant visual ambiguity where multiple tasks share the same scene. This result empirically validates that our method effectively mitigates the vision shortcut, enabling the model to resolve ambiguity through robust instruction following.
To further substantiate this claim, we conducted a quantitative analysis of the conditional entropy $H(\ell\mid v)$ on the LIBERO Goal dataset. We approximated this metric by computing the Negative Log-Likelihood (NLL) of the ground-truth instructions given the visual observations across 40,000 samples. The results are summarized in Table 3.
Table 3: Quantitative analysis of conditional entropy on LIBERO Goal. We report the Negative Log-Likelihood (NLL) and Perplexity (PPL) of instructions given visual observations, serving as a proxy for $H(\ell\mid v)$. Higher NLL/PPL indicates that the model preserves the necessary uncertainty about the task given only vision, preventing the information collapse observed in baselines.
Method NLL (nats/token) $\uparrow$ PPL ($\exp(\text{NLL})$) $\uparrow$ Std Dev (NLL)
QwenGR00T (Baseline) 8.51 4964.1 0.55
LangForce 9.47 12964.9 0.54
LangForce achieves a significantly higher NLL (9.47 nats/token) compared to the QwenGR00T baseline (8.51 nats/token). This difference is even more pronounced in terms of Perplexity (PPL), where LangForce reaches 12964.9 compared to the baseline’s 4964.1. It is worth noting that the language structure in LIBERO Goal is highly repetitive (e.g., “put the [object] in/on the [receptacle]”), which naturally encourages the model to fit these syntactic patterns. Despite this, our method achieves a higher NLL and PPL, indicating that the uncertainty stems primarily from the task-specific nouns (objects and receptacles) rather than the sentence structure. In the context of LIBERO Goal, where visual scenes are inherently ambiguous, the baseline’s lower NLL implies it is “overconfident” in predicting these key nouns solely from vision. This confirms that standard training leads to a collapse in conditional entropy, where the model learns spurious correlations ($v\to\ell$) and ignores the actual need for language disambiguation. In contrast, the higher NLL observed in LangForce indicates that our model prevents the pathological collapse of $H(\ell\mid v)$ observed in standard training, preserving uncertainty levels that align with the inherent ambiguity of visual scenes. By maintaining this necessary uncertainty, our method forces the policy to actively utilize the provided language instruction to resolve ambiguity, directly contributing to the superior performance on the Goal suite.